OpenStudy (anonymous):

How much oxygen, both the moles and grams, is needed to burn all 5000 moles of methane

OpenStudy (anonymous):

CH4+2 O2->CO2+2 H2O If you need to burn 1mole of CH4 you need 2 moles of O2 so for 5000moles we need 10000moles=320000g=320 kg.

OpenStudy (anonymous):

CH4 + 2O2-----------------> CO2 +2H2O mole ratio CH4: O2 = 1: 2 moles of O2 required to burn 5000moles of CH4 = 5000*2 =10000moles mass = moles * Rmm = 10000 * 32 = 320000g